OpenStudy (maxdrizner):

Write the equation of the line that is parallel to the line 4x - 3y = -12 and passes through the point (-3, 4).

OpenStudy (anonymous):

first we have to find the slope of the line given. To do this, we will put it in y = mx + b form and m is the slope. 4x - 3y = -12 (subtract 4x from both sides) -3y = -4x - 12 (now divide both sides by -3) y = -4/-3x - 12/-3 y = 4/3x + 4 slope is 4/3. Parallel lines have the same slope. Now we will use y = mx + b slope(m) = 4/3 (-3,4) x = -3 and y = 4 now lets sub... y = mx + b 4 = 4/3(-3) + b 4 = - 12/3 + b 4 = - 4 + b 4 + 4 = b 8 = b your parallel line is : y = 4/3x + 8